OpenStudy (anonymous):

The number of dogs in a pet store decreased from 35 to 31. What is the percent of decrease in the number of dogs? Round to the nearest percent. A. 11% B. 13% C. 40% D. 97%

OpenStudy (anonymous):

35-31/35 and you get 11% hope this helped :D

OpenStudy (anonymous):

(35-31)/35 =4/35 for % =4*100/35=4*20/7=80/7=11%